Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Message ID: 237     Entry time: Wed Mar 6 10:54:33 2019
Author: Drive Team- Ino&Thierry&Armand 
Type: Telescope Operations 
Category: General 
Subject: DRIVE : High Speed tests 
Status:  
  • The FW issue is not solved . No movement can be done
  • The FW is going to be restarted ==> It takes several hours to put it down and up again
  • Drive movement are stucked 
ELOG V3.1.5-fc6679b